# Crystal Chute Moves on Pipe Dreams

### v1

@@ 1.1. Schubert and key polynomials
Given any permutation $w\in S_{n}$, the Schubert polynomial $\mathfrak{S}_{w}\in{\mathbb{Z}}[x_{1},\dots,x_{n}]$ can be calculated recursively using a sequence of divided difference operators, by the original definition of Lascoux and Schützenberger [LS1], inspired by the work of Demazure [Dem1] and ${}^{\prime}$${}^{\prime}$fand [BGG]. Based on a conjecture of Stanley, the first combinatorial formula for Schubert polynomials was given by Billey, Jockusch, and Stanley using the language of rc-graphs [BJS], with an alternate proof by Fomin and Stanley [FS]. An equivalent combinatorial description for Schubert polynomials was later provided by Fomin and Kirillov [FK], rebranded by Knutson and Miller as reduced pipe dreams [KM], following the conventions of Bergeron and Billey [BB]. Besides being attractive ways to visually represent Schubert polynomials, pipe dreams generalize to flag manifolds the role of the semistandard Young tableaux for Grassmannians, while admitting generalizations to other cohomological contexts.
@@ 2.2. Crystal chute moves
$f_{i}(D)_{+}=\left\{D_{+}\backslash(i,j)\right\}\cup\{(i+1,j-m)\}.$
@@
We claim that $D=f_{i}(D^{\prime})$, where $D_{+}^{\prime}=\left\{D_{+}\backslash(i+1,\ell)\right\}\cup\{(i,n)\}$. By construction, all pipes still enter and exit from the same rows and columns in both $D$ and $D^{\prime}$. As the number of crosses in $D^{\prime}$ is the same as in $D$, then $D^{\prime}\in RP(w)$. Now run the pairing process on row $i$ of $D^{\prime}$. Because $(i+1,\ell)$ was the rightmost unpaired cross in row $i+1$ of $D$, then the cross $(i,n)\in D^{\prime}_{+}$ is the leftmost unpaired cross in row $i$ of $D^{\prime}$. Further note that $(i,\ell)\notin D^{\prime}_{+}$ where $\ell<n$, and so $f_{i}(D^{\prime})\neq 0$. By the argument in the previous paragraph, we see that $m=n-\ell$ satisfies properties (a) and (b) of Definition 2.5 when $j=n$. Therefore, $f_{i}(D^{\prime})=\{D^{\prime}_{+}\backslash(i,n)\}\cup\{(i+1,\ell)\}=D$, and so $D$ is indeed the result of a crystal chute move.
Conversely, suppose that $D=f_{i}(D^{\prime})\neq 0$ for some $D^{\prime}\in RP(w)$. After running the pairing process on row $i$ of $D^{\prime}$, let $(i,j)\in D^{\prime}_{+}$ denote the leftmost unpaired cross in row $i$ of $D^{\prime}$, so that $D_{+}=\{D^{\prime}_{+}\backslash(i,j)\}\cup\{(i+1,j-m)\}$ for $m\in{\mathbb{N}}$ satisfying (a) and (b) of Definition 2.5. Note that the cross $(i+1,j-m)\in D_{+}$ is the rightmost unpaired cross in row $i+1$ after running the pairing process on row $i$ of $D$, since $(i,j)$ was the leftmost unpaired cross in row $i$ of $D^{\prime}$, and the tiles in columns strictly between form a rectangle of crosses by property (b) of $f_{i}(D^{\prime})$. Therefore, $D$ has an unpaired cross in row $i+1$, completing the proof. ∎
@@
Let $n>\ell$ be minimal such that $(i+1,n)\notin D_{+}$. Define a new diagram $e_{i}(D)$ by
$e_{i}(D)_{+}=\left\{D_{+}\backslash(i+1,\ell)\right\}\cup\{(i,n)\}.$
@@
In case there are no unpaired crosses in row $i+1$ after running the pairing process on row $i$ of $D$, then $e_{i}(D)=0$. Otherwise, Lemma 2.9 says that $D$ is the image of the crystal chute move $f_{i}$. Moreover, the proof of Lemma 2.9 shows that the preimage $f_{i}^{-1}(D)$ is precisely the pipe dream $e_{i}(D)_{+}=\left\{D_{+}\backslash(i+1,\ell)\right\}\cup\{(i,n)\},$ and so $e_{i}(D)\in RP(w)\cup\{0\}$ is well-defined. Since $\operatorname{wt}(f_{i}(D))=\operatorname{wt}(D)-\alpha_{i}$ by Proposition 2.8 and $e_{i}(D)=f_{i}^{-1}(D)$ by Lemma 2.9, then $\operatorname{wt}(e_{i}(D))=\operatorname{wt}(D)+\alpha_{i}$. The inverse relationship of $e_{i}$ and $f_{i}$ follows directly from Lemma 2.9 and Definition 2.10. ∎
@@
Let $D\in RP(w)$ be a highest weight pipe dream. Let $c=(i,j_{\ell}^{i})\in D_{+}$ denote the position of $\ell^{\text{th}}$ cross from the left in row $i$ of $D$. Then for any fixed $1\leq\ell\leq\operatorname{wt}(D)_{1}$, the column sequence $(j_{\ell}^{i})$ is weakly increasing.
@@ 3.1. Compatible sequences
[…]
@@ 4.3. Crystal operators on RFCs
We illustrate the pairing process on $r=(\;)(\;)(4)(134)\in RFC(w)$ for $w=[21543]$. Fix $i=1$ and identify $a=4$ as the largest letter in block $(r^{1})=(134)$. We now identify the smallest unpaired letter $b=4$ in block $(r^{2})=(4)$. Since $a=4\not<4=b$, then $a=4$ is unpaired, and we color $a$ red in $r=(\;)(\;)(4)(13{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\pgfsys@color@rgb@stroke{1}{0}{0}\pgfsys@color@rgb@fill{1}{0}{0}4})$.
In step (2), we identify $a^{\prime}=3$ as the letter in block $(r^{1})=(13{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\pgfsys@color@rgb@stroke{1}{0}{0}\pgfsys@color@rgb@fill{1}{0}{0}4})$ immediately to the left of $a=4$. We thus return to step (1) applied to $a=3$ instead. Here again the smallest unpaired letter in block $(r^{2})=(4)$ is $b=4$, which now satisfies $a=3<4=b$. Therefore, these letters get paired in step (1a), as we indicate by coloring them green in $r=(\;)(\;)({\color[rgb]{0.2,0.8,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.8,0.2}\pgfsys@color@rgb@stroke{0.2}{0.8}{0.2}\pgfsys@color@rgb@fill{0.2}{0.8}{0.2}4})(1{\color[rgb]{0.2,0.8,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.8,0.2}\pgfsys@color@rgb@stroke{0.2}{0.8}{0.2}\pgfsys@color@rgb@fill{0.2}{0.8}{0.2}3}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\pgfsys@color@rgb@stroke{1}{0}{0}\pgfsys@color@rgb@fill{1}{0}{0}4})$.
Finally, we return to step (2) to identify the only remaining letter in $(r^{1})=(1{\color[rgb]{0.2,0.8,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.8,0.2}\pgfsys@color@rgb@stroke{0.2}{0.8}{0.2}\pgfsys@color@rgb@fill{0.2}{0.8}{0.2}3}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\pgfsys@color@rgb@stroke{1}{0}{0}\pgfsys@color@rgb@fill{1}{0}{0}4})$ as $a^{\prime}=1$ and reset $a=1$. As there are no unpaired letters in block $(r^{2})=({\color[rgb]{0.2,0.8,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.8,0.2}\pgfsys@color@rgb@stroke{0.2}{0.8}{0.2}\pgfsys@color@rgb@fill{0.2}{0.8}{0.2}4})$, then $a$ remains unpaired in $r=(\;)(\;)({\color[rgb]{0.2,0.8,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.8,0.2}\pgfsys@color@rgb@stroke{0.2}{0.8}{0.2}\pgfsys@color@rgb@fill{0.2}{0.8}{0.2}4})({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\pgfsys@color@rgb@stroke{1}{0}{0}\pgfsys@color@rgb@fill{1}{0}{0}1}{\color[rgb]{0.2,0.8,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.8,0.2}\pgfsys@color@rgb@stroke{0.2}{0.8}{0.2}\pgfsys@color@rgb@fill{0.2}{0.8}{0.2}3}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\pgfsys@color@rgb@stroke{1}{0}{0}\pgfsys@color@rgb@fill{1}{0}{0}4})$. The pairing process on block $i=1$ is now complete, having analyzed all letters in $(r^{1})$.
@@
Consider $r=(\;)(4)(3)(14)\in RFC(w)$ for $w=[21543]$. The outcome of the pairing process on block $i=1$ is $r=(\;)(4)({\color[rgb]{0.2,0.8,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.8,0.2}\pgfsys@color@rgb@stroke{0.2}{0.8}{0.2}\pgfsys@color@rgb@fill{0.2}{0.8}{0.2}3})({\color[rgb]{0.2,0.8,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.8,0.2}\pgfsys@color@rgb@stroke{0.2}{0.8}{0.2}\pgfsys@color@rgb@fill{0.2}{0.8}{0.2}1}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\pgfsys@color@rgb@stroke{1}{0}{0}\pgfsys@color@rgb@fill{1}{0}{0}4})$, where green (resp. red) indicates paired (resp. unpaired) letters. Since there are no unpaired letters in block $(r^{2})=({\color[rgb]{0.2,0.8,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.8,0.2}\pgfsys@color@rgb@stroke{0.2}{0.8}{0.2}\pgfsys@color@rgb@fill{0.2}{0.8}{0.2}3})$, then $e_{1}(r)=0$.
@@ 5.2. Intertwining RFCs and pipe dreams
Now let $D\in RP(w)$, and fix any index $1\leq i<n$ such that $f_{i}(D)\neq 0$. In the notation of Definition 2.5, $f_{i}(D)_{+}=\{D_{+}\backslash(i,j)\}\cup\{(i+1,j-m)\}$, where $(i,j)\in D_{+}$ is the leftmost unpaired cross in row $i$ after running the pairing process on row $i$ of $D$.
@@ 6.1. Truncating permutation from a highest weight pipe dream
        The cross is $\ell^{\text{th}}$ cross from the left in its new row; and
@@
      Repeat steps (a) and (b) until all rows with unfixed crosses have been considered.
@@
[Uncaptioned image] Figure 6.1. A highest weight pipe dream $D$ for the permutation $w=[4726315]\in S_{7}$.
[Uncaptioned image] Figure 6.2. The result of applying steps (1) and (2) of Algorithm 6.1 to $D$; fixed crosses are red.
@@
[Uncaptioned image] Figure 6.3. The diagram after completing the first iteration of step (3); fixed crossed are red.
[Uncaptioned image] Figure 6.4. The diagram $\widetilde{D}$ after completing Algorithm 6.1, with resulting $\operatorname{wt}(\widetilde{D})=(3,5,1,3,1,0)$.
@@
Let $D\in RP(w)$ be a highest weight pipe dream. Let $c=(i,j^{i}_{\ell})\in D_{+}$ denote $\ell^{\text{th}}$ cross from the left in row $i$ of $D$. Then for any fixed $1\leq\ell\leq\operatorname{wt}(D)_{1}$, the shifted column sequence $(j^{i}_{\ell}+i-1)$ is strictly increasing.
@@
Let $D\in RP(w)$ be a highest weight pipe dream. Throughout the proof, given a cross $c$, denote by $\tilde{c}$ the image of $c$ after it is fixed by the algorithm, and denote by $c^{\prime}$ the closest cross to the left of $c$ in its original row. Denote the row index of $c$ recorded from the top down as $\operatorname{row}(c)$.
@@
We now verify the base case $\operatorname{row}(\tilde{c}_{2})<\operatorname{row}(\tilde{c}_{1})$ for the second claim. Recall that $c_{1}$ has already been fixed to obtain $\tilde{c}_{1}$. By the same argument, any cross below $c_{2}$ and in the same column must lie in a row strictly below the fixed cross $\tilde{c}^{\prime}_{2}$, so that $c_{2}$ can move freely down its column, at least as far down as to join $\tilde{c}^{\prime}_{2}$ to become the second cross in its new row. Moving $c_{2}$ in this manner clearly satisfies (i) and (ii). Denote by $\operatorname{row}(\tilde{c}_{2})=m$ and $\operatorname{row}(\tilde{c}_{1})=n$, in which case there are crosses on the diagonal in positions $(m,m)$ and $(n,n)$ from step (2) and property (ii). If $n<m$, then in fact $c_{1}$ could have moved farther down into row $m$ without moving through any crosses, based on the relative row and column ordering of crosses with iteration index $\ell=1$, and the fact that $c_{1}$ is the lowest cross with index $\ell=2$. Since $\tilde{c}_{1}$ moves as low as possible by the algorithm, then $\operatorname{row}(\tilde{c}_{2})=m\leq n=\operatorname{row}(\tilde{c}_{1})$. Since both $c_{1}$ and $c_{2}$ are free to move down their respective columns, at least as far as to rejoin $\tilde{c}^{\prime}_{1}$ and $\tilde{c}^{\prime}_{2}$ which satisfy $\operatorname{row}(\tilde{c}^{\prime}_{2})<\operatorname{row}(\tilde{c}^{\prime}_{1})$, then additionally $m\neq n$. Therefore, $\operatorname{row}(\tilde{c}_{2})<\operatorname{row}(\tilde{c}_{1})$, confirming the base case for this inequality.
For the inductive step, assume that crosses $c_{1}$ through $c_{k}$ have been successfully placed by the algorithm. Further assume that $\operatorname{row}(\tilde{c}_{j})<\operatorname{row}(\tilde{c}_{j-1})$ for all $1\leq j\leq k$ such that $c_{j-1}$ and $c_{j}$ have the same iteration index. There are two cases to consider in the inductive step: either crosses $c_{k}$ and $c_{k+1}$ have the same iteration index or they do not. If $c_{k}$ has iteration index $\ell-1$ and $c_{k+1}$ has iteration index $\ell$, then the same argument supplied in the base case, instead applying Lemma 6.3 to the $(\ell-1)^{\text{st}}$ column sequence, shows that the first cross having iteration index $\ell$, namely $c_{k+1}$, can be placed in a manner satisfying (i), (ii), and (iii). The claim concerning the row indices does not apply in case $c_{k}$ and $c_{k+1}$ have different iteration indices.
@@
6.2. Weak Edelman-Greene insertion
To prove Theorem 6.5, we relate $\pi_{D}$ with the sorting permutation $\pi_{r}$ from Theorem 4.10. To obtain $\pi_{r}$ following [AS], we must apply a modification of Edelman-Greene insertion. We thus continue by reviewing this insertion algorithm.

### v2

@@ 1.1. Schubert and key polynomials
Given any permutation $w\in S_{n}$, the Schubert polynomial $\mathfrak{S}_{w}\in{\mathbb{Z}}[x_{1},\dots,x_{n}]$ can be calculated recursively using a sequence of divided difference operators, by the original definition of Lascoux and Schützenberger [LS1], inspired by the work of Demazure [Dem1] and Bernstein–Gelfand–Gelfand [BGG]. Based on a conjecture of Stanley, the first combinatorial formula for Schubert polynomials was given by Billey, Jockusch, and Stanley using the language of rc-graphs [BJS], with an alternate proof by Fomin and Stanley [FS]. An equivalent combinatorial description for Schubert polynomials was later provided by Fomin and Kirillov [FK], rebranded by Knutson and Miller as reduced pipe dreams [KM], following the conventions of Bergeron and Billey [BB]. Besides being attractive ways to visually represent Schubert polynomials, pipe dreams generalize to flag manifolds the role of the semistandard Young tableaux for Grassmannians, while admitting generalizations to other cohomological contexts.
@@ 2.2. Crystal chute moves
$f_{i}(D)_{+}=D_{+}\backslash\left\{(i,j)\right\}\cup\{(i+1,j-m)\}.$
@@
We claim that $D=f_{i}(D^{\prime})$, where $D_{+}^{\prime}=D_{+}\backslash\left\{(i+1,\ell)\right\}\cup\{(i,n)\}$. By construction, all pipes still enter and exit from the same rows and columns in both $D$ and $D^{\prime}$. As the number of crosses in $D^{\prime}$ is the same as in $D$, then $D^{\prime}\in RP(w)$. Now run the pairing process on row $i$ of $D^{\prime}$. Because $(i+1,\ell)$ was the rightmost unpaired cross in row $i+1$ of $D$, then the cross $(i,n)\in D^{\prime}_{+}$ is the leftmost unpaired cross in row $i$ of $D^{\prime}$. Further note that $(i,\ell)\notin D^{\prime}_{+}$ where $\ell<n$, and so $f_{i}(D^{\prime})\neq 0$. By the argument in the previous paragraph, we see that $m=n-\ell$ satisfies properties (a) and (b) of Definition 2.5 when $j=n$. Therefore, $f_{i}(D^{\prime})=D^{\prime}_{+}\backslash\{(i,n)\}\cup\{(i+1,\ell)\}=D$, and so $D$ is indeed the result of a crystal chute move.
Conversely, suppose that $D=f_{i}(D^{\prime})\neq 0$ for some $D^{\prime}\in RP(w)$. After running the pairing process on row $i$ of $D^{\prime}$, let $(i,j)\in D^{\prime}_{+}$ denote the leftmost unpaired cross in row $i$ of $D^{\prime}$, so that $D_{+}=D^{\prime}_{+}\backslash\{(i,j)\}\cup\{(i+1,j-m)\}$ for $m\in{\mathbb{N}}$ satisfying (a) and (b) of Definition 2.5. Note that the cross $(i+1,j-m)\in D_{+}$ is the rightmost unpaired cross in row $i+1$ after running the pairing process on row $i$ of $D$, since $(i,j)$ was the leftmost unpaired cross in row $i$ of $D^{\prime}$, and the tiles in columns strictly between form a rectangle of crosses by property (b) of $f_{i}(D^{\prime})$. Therefore, $D$ has an unpaired cross in row $i+1$, completing the proof. ∎
@@
Let $q>\ell$ be minimal such that $(i+1,q)\notin D_{+}$. Define a new diagram $e_{i}(D)$ by
$e_{i}(D)_{+}=D_{+}\backslash\left\{(i+1,\ell)\right\}\cup\{(i,q)\}.$
@@
In case there are no unpaired crosses in row $i+1$ after running the pairing process on row $i$ of $D$, then $e_{i}(D)=0$. Otherwise, Lemma 2.9 says that $D$ is the image of the crystal chute move $f_{i}$. Moreover, the proof of Lemma 2.9 shows that the preimage $f_{i}^{-1}(D)$ is precisely the pipe dream $e_{i}(D)_{+}=D_{+}\backslash\left\{(i+1,\ell)\right\}\cup\{(i,q)\},$ and so $e_{i}(D)\in RP(w)\cup\{0\}$ is well-defined. Since $\operatorname{wt}(f_{i}(D))=\operatorname{wt}(D)-\alpha_{i}$ by Proposition 2.8 and $e_{i}(D)=f_{i}^{-1}(D)$ by Lemma 2.9, then $\operatorname{wt}(e_{i}(D))=\operatorname{wt}(D)+\alpha_{i}$. The inverse relationship of $e_{i}$ and $f_{i}$ follows directly from Lemma 2.9 and Definition 2.10. ∎
@@
Let $D\in RP(w)$ be a highest weight pipe dream. Let $c=(i,j_{\ell}^{i})\in D_{+}$ denote the position of the $\ell^{\text{th}}$ cross from the left in row $i$ of $D$. Then for any fixed $1\leq\ell\leq\operatorname{wt}(D)_{1}$, the column sequence $(j_{\ell}^{i})$ is weakly increasing.
@@ 3.1. Compatible sequences
$D_{+}=\{(i_{1},j_{1}),\dots,(i_{p},j_{p})\}$ $\ \longmapsto\ \beta=(i_{1},\dots,i_{p})\in C(\mathbf{a})\ \ \text{for}\ \ \mathbf{a}=(i_{1}+j_{1}-1,\dots,i_{p}+j_{p}-1),$
@@ 4.3. Crystal operators on RFCs
We illustrate the pairing process on $r=(\;)(\;)(4)(134)\in RFC(w)$ for $w=[21543]$. Fix $i=1$ and identify $a=4$ as the largest letter in block $(r^{1})=(134)$. We now identify the smallest unpaired letter $b=4$ in block $(r^{2})=(4)$. Since $a=4\not<4=b$, then $a=4$ is unpaired, and we color $a$ red in $r=(\;)(\;)(4)(13{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}4})$.
In step (2), we identify $a^{\prime}=3$ as the letter in block $(r^{1})=(13{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}4})$ immediately to the left of $a=4$. We thus return to step (1) applied to $a=3$ instead. Here again the smallest unpaired letter in block $(r^{2})=(4)$ is $b=4$, which now satisfies $a=3<4=b$. Therefore, these letters get paired in step (1a), as we indicate by coloring them green in $r=(\;)(\;)({\color[rgb]{0.2,0.8,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.8,0.2}4})(1{\color[rgb]{0.2,0.8,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.8,0.2}3}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}4})$.
Finally, we return to step (2) to identify the only remaining letter in $(r^{1})=(1{\color[rgb]{0.2,0.8,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.8,0.2}3}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}4})$ as $a^{\prime}=1$ and reset $a=1$. As there are no unpaired letters in block $(r^{2})=({\color[rgb]{0.2,0.8,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.8,0.2}4})$, then $a$ remains unpaired in $r=(\;)(\;)({\color[rgb]{0.2,0.8,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.8,0.2}4})({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}1}{\color[rgb]{0.2,0.8,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.8,0.2}3}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}4})$. The pairing process on block $i=1$ is now complete, having analyzed all letters in $(r^{1})$.
@@
Consider $r=(\;)(4)(3)(14)\in RFC(w)$ for $w=[21543]$. The outcome of the pairing process on block $i=1$ is $r=(\;)(4)({\color[rgb]{0.2,0.8,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.8,0.2}3})({\color[rgb]{0.2,0.8,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.8,0.2}1}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}4})$, where green (resp. red) indicates paired (resp. unpaired) letters. Since there are no unpaired letters in block $(r^{2})=({\color[rgb]{0.2,0.8,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.8,0.2}3})$, then $e_{1}(r)=0$.
@@ 5.2. Intertwining RFCs and pipe dreams
Now let $D\in RP(w)$, and fix any index $1\leq i<n$ such that $f_{i}(D)\neq 0$. In the notation of Definition 2.5, $f_{i}(D)_{+}=D_{+}\backslash\{(i,j)\}\cup\{(i+1,j-m)\}$, where $(i,j)\in D_{+}$ is the leftmost unpaired cross in row $i$ after running the pairing process on row $i$ of $D$.
@@ 6.1. Truncating permutation from a highest weight pipe dream
        The cross is the $\ell^{\text{th}}$ cross from the left in its new row; and
@@
      Repeat steps (a) and (b) once within each row until all rows with unfixed crosses have been considered.
@@
Figure 6.1. A highest weight pipe dream $D$ for the permutation $w=[4726315]\in S_{7}$.
Figure 6.2. The result of applying steps (1) and (2) of Algorithm 6.1 to $D$; fixed crosses are red.
@@
Figure 6.3. The diagram after completing the first iteration of step (3); fixed crossed are red.
Figure 6.4. The diagram $\widetilde{D}$ after completing Algorithm 6.1, with resulting $\operatorname{wt}(\widetilde{D})=(3,5,1,3,1,0)$.
@@
Let $D\in RP(w)$ be a highest weight pipe dream. Let $c=(i,j^{i}_{\ell})\in D_{+}$ denote the $\ell^{\text{th}}$ cross from the left in row $i$ of $D$. Then for any fixed $1\leq\ell\leq\operatorname{wt}(D)_{1}$, the shifted column sequence $(j^{i}_{\ell}+i-1)$ is strictly increasing.
@@
Let $D\in RP(w)$ be a highest weight pipe dream. Throughout the proof, given a cross $c$, denote by $\tilde{c}$ the image of $c$ after it is fixed by the algorithm, and denote by $c^{\prime}$ the closest cross to the left of $c$ in its original row. Denote by $\tilde{c}^{\prime}=(\tilde{c})^{\prime}$ the cross to the immediate left of $\tilde{c}$ in its original row. Denote the row index of $c$ recorded from the top down as $\operatorname{row}(c)$.
@@
We now verify the base case $\operatorname{row}(\tilde{c}_{2})<\operatorname{row}(\tilde{c}_{1})$ for the second claim. Recall that $c_{1}$ has already been fixed to obtain $\tilde{c}_{1}$. By the same argument, any cross below $c_{2}$ and in the same column must lie in a row strictly below the fixed cross $\tilde{c}^{\prime}_{2}$, so that $c_{2}$ can move freely down its column, at least as far down as to join $\tilde{c}^{\prime}_{2}$ to become the second cross in its new row. Moving $c_{2}$ in this manner clearly satisfies (i) and (ii). Denote by $\operatorname{row}(\tilde{c}_{2})=m$ and $\operatorname{row}(\tilde{c}_{1})=q$, in which case there are crosses on the diagonal in positions $(m,m)$ and $(q,q)$ from step (2) and property (ii). If $q<m$, then in fact $c_{1}$ could have moved farther down into row $m$ without moving through any crosses, based on the relative row and column ordering of crosses with iteration index $\ell=1$, and the fact that $c_{1}$ is the lowest cross with index $\ell=2$. Since $\tilde{c}_{1}$ moves as low as possible by the algorithm, then $\operatorname{row}(\tilde{c}_{2})=m\leq q=\operatorname{row}(\tilde{c}_{1})$. In order that $\tilde{c}_{2}$ satisfies (ii), then additionally $m\neq q$. Therefore, $\operatorname{row}(\tilde{c}_{2})<\operatorname{row}(\tilde{c}_{1})$, confirming the base case for this inequality.
For the inductive step, assume that crosses $c_{1}$ through $c_{k}$ have been successfully placed by the algorithm. Further assume that $\operatorname{row}(\tilde{c}_{j})<\operatorname{row}(\tilde{c}_{j-1})$ for all $2\leq j\leq k$ such that $c_{j-1}$ and $c_{j}$ have the same iteration index. There are two cases to consider in the inductive step: either crosses $c_{k}$ and $c_{k+1}$ have the same iteration index or they do not. If $c_{k}$ has iteration index $\ell-1$ and $c_{k+1}$ has iteration index $\ell$, then the same argument supplied in the base case, instead applying Lemma 6.3 to the $(\ell-1)^{\text{st}}$ column sequence, shows that the first cross having iteration index $\ell$, namely $c_{k+1}$, can be placed in a manner satisfying (i), (ii), and (iii). The claim concerning the row indices does not apply in case $c_{k}$ and $c_{k+1}$ have different iteration indices.
@@
6.2. Weak Edelman–Greene insertion
To prove Theorem 6.5, we relate $\pi_{D}$ with the sorting permutation $\pi_{r}$ from Theorem 4.10. To obtain $\pi_{r}$ following [AS], we must apply a modification of Edelman–Greene insertion. We thus continue by reviewing this insertion algorithm.